# Data-centric Prediction Explanation via Kernelized Stein Discrepancy

### v1

@@ Abstract
Existing example-based prediction explanation methods often bridge test and training data points through the model’s parameters or latent representations. While these methods offer clues to the causes of model predictions, they often exhibit innate shortcomings, such as incurring significant computational overhead or producing coarse-grained explanations. This paper presents a Highly-precise and Data-centric Explanation (HD-Explain), a straightforward prediction explanation method exploiting properties of Kernelized Stein Discrepancy (KSD). Specifically, the KSD uniquely defines a parameterized kernel function for a trained model that encodes model-dependent data correlation. By leveraging the kernel function, one can identify training samples that provide the best predictive support to a test point efficiently. We conducted thorough analyses and experiments across multiple classification domains, where we show that HD-Explain outperforms existing methods from various aspects, including 1) preciseness (fine-grained explanation), 2) consistency, and 3) computation efficiency, leading to a surprisingly simple, effective, and robust prediction explanation solution.
Machine Learning, ICML
@@ 1 Introduction
This paper presents Highly-precise and Data-centric Explanation (HD-Explain), a post-hoc, model-aware, example-based explanation solution for neural classifiers. Instead of relying on data co-influence on model parameters or feature representation similarity, HD-Explain retains the influence chain between training and test data points by exploiting the underrated properties of Kernelized Stein Discrepancy (KSD) between the trained predictive model and its training dataset. Specifically, we note that the Stein operator augmented kernel uniquely defines a pairwise data correlation (in the context of a trained model) whose expectation on the training dataset results in the minimum KSD (as a discrete approximation) compared to that of the dataset sampled from different distributions. By exploiting this property, we can 1) reveal a subset of training data points that provides the best predictive support to the test point and 2) identify the potential distribution mismatch among training data points. Jointly leveraging these advantages, HD-Explain can produce explanations faithful to the original trained model with a corresponding confidence level.
We conducted a wide range of qualitative and quantitative experiments and analyses to demonstrate the effectiveness and efficiency of HD-Explain. Our empirical results show that the HD-Explain offers fine-grained, instance-level explanations with remarkable computational efficiency on multiple applications, compared to well-known example-based prediction explanation solutions. Given its algorithmic simplicity and scalability, we believe HD-Explain is an excellent addition to the literature on example-based prediction explanation research.
@@ 2.1 Post-hoc Classifier Explanation by Examples
Post-hoc Classifier Explanation by Examples (a.k.a prototypes, representers) refers to a category of classifier explanation approaches that pick a subset of training data points as prediction explanations without accessing the model training process. Its research history spans from the model-intrinsic approach to the recent impact-based approach.
Model-inherent approaches are essentially a feature of self-explanatory machine learning models such as $k$-nearest neighbor (Peterson, 2009) or decision tree; For a given test data point, similar data points on the raw feature space can be efficiently selected as explanations through the inherent decision making mechanism of the self-explanatory machine learning models. In fact, attracted by their inherent explanatory power, multiple well-known works attempted to compile complex black-box models into self-explanatory models for enabling prediction explanation (Frosst & Hinton, 2017), while computationally inefficient.
To unlock the general explanatory power applicable to black-box models, multiple later studies suggest to fall back to statistics-based solutions, looking for prototype samples that could represent common data points or play critical roles in data distribution. MMD-critic (Kim et al., 2016) and Normative and Comprarative explanations (Cai et al., 2019) are the well-known examples in this category. Unfortunately, those approaches often overlook the characteristics of trained models, making their prediction explanations general to the training dataset rather than a trained model instance.
@@ 2.2 Kernelized Stein Discrepancy
$\mathbb{E}_{x\sim p}[\phi(x)\nabla_{x}\log p(x)+\nabla_{x}\phi(x)]=0,\quad\forall f.$
The above expression could be further abstracted to use function operator $\mathcal{A}_{p}$ (a.k.a Stein operator) such that
@@
where $\kappa_{p}(x,x^{\prime})=\mathcal{A}_{p}^{x}\mathcal{A}_{p}^{x^{\prime}}k(x,x^{\prime})$ that can work with arbitrary kernel function $k(x,x^{\prime})$. See Appendix B for expanded derivations.
@@
3 Example-based Prediction Explanation via Kernelized Stein Discrepancy (HD-Explain)
Consider a trained classifier $f_{\theta}$ as the outcome of a training process with Maximum Likelihood Estimation (MLE)
@@ 3 Example-based Prediction Explanation via Kernelized Stein Discrepancy (HD-Explain)
$=-\underbrace{\mathbb{E}_{(\mathbf{x},y)\sim P_{D}}[\log P_{\theta}(y|\mathbf{x})]}_{\text{Likelihood}}+\underbrace{\mathbb{E}_{(\mathbf{x},y)\sim P_{D}}[\log P_{D}(y|\mathbf{x})]}_{\text{constant}},$
which, in turn, is proven recently to align with minimizing Stein Discrepancy in the form of gradient descent (Liu & Wang, 2016)
$\nabla_{\theta}\mathbb{D}_{\text{KL}}(P_{D},P_{\theta})=-\mathbb{S}(P_{D},P_{\theta}).$
The chain of reasoning above shows that a well-trained classifier $f_{\theta}$ through gradient-descent should lead to minimum Stein Discrepancy between the training dataset distribution and the model encoded distribution $\mathbb{S}(P_{D},P_{\theta})$  [^1] Since $P_{D}$ is discrete distribution while $P_{\theta}$ is continuous, the Stein Discrepancy between the two distributions will not recap Stein Identity ($=0$) with a limited number of training data points.. We can empirically verify the connection through simple examples as shown in Figure 1, where the changes in training data distribution would result in larger KSD compared to that of the original training data distribution. Intuitively, the connection shows that there is a tie between a model and its training data points, encoded in the form of a stein kernel function $k_{\theta}(\cdot,\cdot)$ defined on each pair of data points. As the kernel function is conditioned on model $f_{\theta}$, we note it is an encoding of data correlation under the context of a trained model, which paves the foundation of the example-based prediction explanation as shown next.
@@ 3.1 KSD between Model and Training Data
where we denote data point $(\mathbf{x}_{a},y_{a})$ with $a$ for clean notation. The only model-dependent factor in the above decomposition is a derivative $\nabla_{\mathbf{x},y}\log P_{\theta}(\mathbf{x},y)$ (for both data $a$ or $b$). Hence, to utilize KSD for prediction explanation, we first need to unravel the derivative of the trained model with respect to a whole data point (including its label). Indeed, while the gradient of a classifier $\nabla_{\mathbf{x}}\log P_{\theta}(y|\mathbf{x})$ on its input feature $\mathbf{x}$ is straightforward, the gradient to the entire data points is less evident due to the lack of definition of $P_{\theta}(\mathbf{x})$.
To obtain the joint distribution of $P_{\theta}(\mathbf{x},y)$ for facilitating KSD computation, we propose to set $P_{\theta}(\mathbf{x})\equiv P_{D}(\mathbf{x})$ as uniform distribution. Although this setting appears hasty, we note $P_{D}$ represents the uniform data distribution in dataset $D$, which, while reflects, but is not the complex unknown distribution from which the data is sampled. In particular, for a data point $(\mathbf{x},y)$ in dataset D with i.i.d assumption, $P_{D}(\mathbf{x})$ is indeed uniformly distributed, whereas $P_{D}(y|\mathbf{x})$ is a delta distribution with probability $1$.
With the above setting, the score function $\nabla_{\mathbf{x},y}\log P_{\theta}(\mathbf{x},y)$ in the Stein operator $\mathcal{A}_{\theta}$ could be derived as a concatenation of the gradient of model $f_{\theta}(\mathbf{x})_{y}$ to its input $\mathbf{x}$ and its probabilistic prediction $f_{\theta}(\mathbf{x})$, since
@@
$=[\nabla_{\mathbf{x}}\log\mathbf{y}^{\top}f_{\theta}(\mathbf{x})||\nabla_{\mathbf{y}}\log\mathbf{y}^{\top}f_{\theta}(\mathbf{x})]+[\mathbf{0}||0]$
$=[\nabla_{\mathbf{x}}\log f_{\theta}(\mathbf{x})_{y}||f_{\theta}(\mathbf{x})],$
where [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes concatenation operation, and we use one-hot encoding $\mathbf{y}$ of $y$ to clarify the gradient computation.
@@
Combining Equation 1 and 2, we can estimate the correlation of any pairs of training data points conditioned on the trained machine learning model. Computationally, since a score function $\nabla_{\mathbf{x},y}\log P_{\theta}(\mathbf{x},y)$ depends on a single data point, its outputs of the training set could be pre-computed and cached for accelerate the kernel computation. In particular, the output dimension of the score function is simply $m+k$ for data with $m$ dimensional features and $k$ class labels. Compared to the existing solutions, whose training data cache (or influence) are bounded by the dimension of model parameters (such as Influence function, TracIn, RPS, RPS-JLE), the explanation method built on KSD would come with a significant advantage in terms of scalability. We note this statement is generally true for neural network based classifiers, whenever the size of model parameters is far larger than the data dimension.
@@
Figure 2: Demonstration of HD-Explain on 2D Rectangular synthetic dataset. Top-left shows the training dataset with three classes. Top-right shows the explanation support of training data points to a given test point (as black cross), where green shows a higher KSD kernel value. The bottom shows the distribution of KSD kernel value to the test point, where only a small number of training data points provide strong support to this prediction.
3.2 Prediction Explanation via KSD
@@
Now, we present how to extend the previously described kernel function to support the prediction explanation. As stated in Equation 1, computing the kernel function requires to access features and labels of a data point. While the ground-truth label information is provided in the training set, it is inaccessible during inference time given a test data point. Hence, to utilize the KSD kernel function, we place predicted class $\hat{y}_{t}$ of the test point $\mathbf{x}_{t}$ as a label to construct a complete data point $(\mathbf{x}_{t},\hat{y}_{t})$.
At this stage, searching supportive training examples for a given test point $\mathbf{x}_{t}$ could be reduced to look for top-k training data points that maximize the KSD defined kernel. For top-1 explanation sample, in particular, we look for
$\operatorname*{argmax}_{i}\kappa_{\theta}([\mathbf{x}_{t}||\hat{\mathbf{y}}_{t}],[\mathbf{x}_{i}||\mathbf{y}_{i}]),$ (3)
where, again, we use one-hot encoding for data labels.
@@
Figure 2 demonstrates a simple example on a 2d synthetic dataset. By iterating over all training data points and evaluating $\kappa_{\theta}(\mathbf{d},\cdot)$, we highlight that only a small number of training data points provide a strong influence on a particular prediction (see Figure  2 (bottom)).
(a) Highly-Confident Correct Prediction
(b) Low-Confident Correct Prediction
(c) Low-Confident Incorrect Prediction
Figure 3: Qualitative explanation comparison among candidate example-based explanation methods. We show three scenarios where the target model makes 1) a highly-confident prediction that matches ground truth label, 2) a low-confident prediction that matches ground truth label, 3) low-confident prediction that does not match ground truth label (which is a bird). For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example.
(a) Overian Cancer Histopathology
(b) Brain Tumor MRI
Figure 4: Qualitative explanation comparison among candidate example-based explanation methods on Overian Cancer histopathology and Brain Tumor MRI datasets. We show two test data points that are predicted to belong to the same class in each dataset. Red triangle in the top right corner of an image shows the duplicate explanations across test samples.
4 Evaluations and Analysis
@@
In this section, we conduct several qualitative and quantitative experiments to demonstrate various properties of the proposed HD-Explain solution and analyze it with existing example-based solutions to understand where and how the approach can be used in practice. Codes are available in the supplementary materials.
@@
Datasets: We are particularly interested in classification tasks that are highly sensitive to trustworthiness and with a clear demand for prediction explanations. To this end, we designed our experiments to include multiple disease classification datasets, where diagnosis explanation is highly desired. We also introduced synthetic and benchmark classification datasets to deliver the main idea without the need for medical background knowledge. Concretely, the primary datasets used in our experiments include CIFAR-10 ($32\times 32\times 3$), Brain Tumor (Magnetic Resonance Imaging, $128\times 128\times 3$), and Ovarian Cancer (Histopathology Images, $128\times 128\times 3$). More details of the datasets are listed in the Appendix.
@@
Baselines: The baseline explainers used in our experiments include Influence Function, Representer Point Selection, and TracIn. While other variants of these baseline explainers exist (Barshan et al., 2020; Sui et al., 2021; Chen et al., 2021), we note they don’t offer fundamental performance improvements over the classic ones. In addition, as Influence Function and TracIn face scalability issues, we limit the influence of parameters to the last layer of the model so that they can work with models that contain a large number of parameters. Our experiments use ResNet-18 as the backbone model architecture (with around $11$ million trainable parameters) for all image datasets. Finally, we also introduce an HD-Explain variant (HD-Explain*) to match the last layer setting of other baseline models, even though HD-Explain can scale up to the whole model without computation pressure (see  Appendix F for details).
@@
Metrics: In existing example-based explanation works, the experimental results are often demonstrated qualitatively, as visualized explanation instances, without quantitative evaluation, which, we believe, is subjective and hardly convincing. To obtain a comprehensive understanding of the effectiveness of explanation methods, we proposed several quantitative metrics in this paper.
@@
  • Hit Rate: Hit rate shows how likely an explanation sample hits the desired example cases. Here, the desired examples are guaranteed to be undisputed. Specifically, we modify a training data point with minor augmentations (adding noise or flipping horizontally) and use it as a test data point, such that the best explanation for the generated test data point should be the original data point in the training set.
@@
  • Coverage: Given n test data points, we show how many unique explanation samples an explanation method can produce when configuring to return top-k training samples. Formally,
@@
    where $e_{i}$ is the set of top-k explanations for test data point $i$. Larger coverage is preferred as it suggests the higher granularity (per test point) of the explanation.
  • Run Time: Measuring the run time of explanation methods. We measure the wall clock time.

### v2

@@ Abstract
Existing example-based prediction explanation methods often bridge test and training data points through the model’s parameters or latent representations. While these methods offer clues to the causes of model predictions, they often exhibit innate shortcomings, such as incurring significant computational overhead or producing coarse-grained explanations. This paper presents a Highly-precise and Data-centric Explanation (HD-Explain) prediction explanation method that exploits properties of Kernelized Stein Discrepancy (KSD). Specifically, the KSD uniquely defines a parameterized kernel function for a trained model that encodes model-dependent data correlation. By leveraging the kernel function, one can identify training samples that provide the best predictive support to a test point efficiently. We conducted thorough analyses and experiments across multiple classification domains, where we show that HD-Explain outperforms existing methods from various aspects, including 1) preciseness (fine-grained explanation), 2) consistency, and 3) computation efficiency, leading to a surprisingly simple, effective, and robust prediction explanation solution.
@@ 1 Introduction
This paper presents Highly-precise and Data-centric Explanation (HD-Explain), a post-hoc, model-aware, example-based explanation solution for neural classifiers. Instead of relying on data co-influence on model parameters or feature representation similarity, HD-Explain retains the influence chain between training and test data points by exploiting the underrated properties of Kernelized Stein Discrepancy (KSD) (Liu et al., 2016) between the trained predictive model and its training dataset. Specifically, we note that the Stein operator augmented kernel uniquely defines a pairwise data correlation (in the context of a trained model) whose expectation on the training dataset results in the minimum KSD (as a discrete approximation) compared to that of the dataset sampled from different distributions. By exploiting this property, we can 1) reveal a subset of training data points that provides the best predictive support to the test point and 2) identify the potential distribution mismatch among training data points. Jointly leveraging these advantages, HD-Explain can produce explanations faithful to the original trained model.
The contributions of our work are summarized as follows: 1) We propose a novel example-based explanation method. 2) We propose several quantitative evaluation metrics to measure the correctness and effectiveness of generated explanations. 3) We perform a thorough evaluation comparing several existing explanation methods across a wide set of classification tasks.
@@
Our findings conclude that HD-Explain offers fine-grained, instance-level explanations with remarkable computational efficiency, compared to well-known example-based prediction explanation methods. Its algorithmic simplicity, effective performance and scalability enables its deployment to real world scenarios where transparency and trustworthiness is essential.
@@ 2.1 Post-hoc Classifier Explanation by Examples
Post-hoc Classifier Explanation by Examples (a.k.a prototypes, representers) refers to a category of classifier explanation approaches that pick a subset of training data points as prediction explanations without accessing the model training process. Its research history spans from the model-intrinsic approach (Molnar, 2020) to the recent impact-based approach (Li et al., 2020).
Model-inherent approaches (Molnar, 2020) refer to machine learning models that are considered interpretable such as $k$-nearest neighbor (Peterson, 2009) or decision tree; For a given test data point, similar data points on the raw feature space can be efficiently selected as explanations through the inherent decision making mechanism of the self-explanatory machine learning models. In fact, attracted by their inherent explanatory power, multiple well-known works attempted to compile complex black-box models into self-explanatory models for enabling prediction explanation (Frosst & Hinton, 2017), while computationally inefficient.
To unlock the general explanatory power applicable to black-box models, multiple later studies suggested to fall back to statistics-based solutions, looking for prototype samples that represent data points that either are common in the dataset or play critical roles in data distribution. MMD-critic (Kim et al., 2016) and Normative and Comparative explanations (Cai et al., 2019) are the well-known examples in this category. Unfortunately, those approaches are often with strong assumption of good prototypes (as of training data distribution) (Li et al., 2020), making their prediction explanations general to the training dataset rather than a trained model instance.
@@ 2.2 Kernelized Stein Discrepancy
$\mathbb{E}_{x\sim p}[\phi(x)\nabla_{x}\log p(x)+\nabla_{x}\phi(x)]=0,\quad\forall\phi.$
The identity can characterize distribution $p(x)$ such that it is often served to assess the goodness-of-fit (Kubokawa, 2024). The above expression could be further abstracted to use function operator $\mathcal{A}_{p}$ (a.k.a Stein operator) such that
@@
where $\kappa_{p}(x,x^{\prime})=\mathcal{A}_{p}^{x}\mathcal{A}_{p}^{x^{\prime}}k(x,x^{\prime})$ that can work with arbitrary kernel function $k(x,x^{\prime})$. See Appendix C for expanded derivations.
@@
3 Highly-precise and Data-centric Explanation
HD-Explain is an example-based prediction explanation method based on Kernelized Stein Discrepancy. Consider a trained classifier $f_{\theta}$ as the outcome of a training process with Maximum Likelihood Estimation (MLE)
@@ 3 Highly-precise and Data-centric Explanation
$=-\underbrace{\mathbb{E}_{(\mathbf{x},y)\sim P_{D}}[\log P_{\theta}(y|\mathbf{x})]}_{\text{Likelihood}}+\underbrace{\mathbb{E}_{(\mathbf{x},y)\sim P_{D}}[\log P_{D}(y|\mathbf{x})]}_{\text{constant}}+\underbrace{\mathbb{E}_{(\mathbf{x},y)\sim P_{D}}[\log P_{D}(\mathbf{x}/P_{\theta}(\mathbf{x})]}_{\text{constant as $\theta$ does not model inputs}},$
which, in turn, is proven to align with minimizing KSD in the form of gradient descent (Liu & Wang, 2016)
$\nabla_{\epsilon}\mathbb{D}_{\text{KL}}(P_{D},P_{\theta})\rvert_{\epsilon}=-\mathbb{S}(P_{D},P_{\theta}),$
where $\epsilon$ is the step size of gradient decent.
The chain of reasoning above shows that a well-trained classifier $f_{\theta}$ through gradient-descent should lead to minimum discrepancy between the training dataset distribution and the model encoded distribution $\mathbb{S}(P_{D},P_{\theta})$  [^1] Since $P_{D}$ is discrete distribution while $P_{\theta}$ is continuous, the Discrepancy between the two distributions will not recap Stein Identity ($=0$) with a limited number of training data points.. We empirically verify the connection through simple examples as shown in Figure 1, where the changes in training data distribution would result in larger KSD compared to that of the original training data distribution. Intuitively, the connection shows that there is a tie between a model and its training data points, encoded in the form of a Stein kernel function $k_{\theta}(\cdot,\cdot)$ defined on each pair of data points. As the kernel function is conditioned on model $f_{\theta}$, we note it is an encoding of data correlation under the context of a trained model, which paves the foundation of the example-based prediction explanation.
@@ 3.1 KSD between Model and Training Data
where we denote data point $(\mathbf{x}_{a},y_{a})$ with $a$ for clean notation. The only model-dependent factor in the above decomposition is a derivative $\nabla_{\mathbf{x},y}\log P_{\theta}(\mathbf{x},y)$ (for both data $a$ or $b$).
However, as the KSD only models the discrepancy between joint distributions rather than conditional distributions, it is challenging to estimate discrepancy between predictive models $P_{\theta}(y|\mathbf{x})$ and its training set $\mathbf{z}=(\mathbf{x},y)\sim P_{D}$ without including $P_{\theta}(\mathbf{x})$ into consideration, even though the marginal distribution $P_{\theta}(\mathbf{x})$ is not estimated by the predictive model at all.
Inspired by the previous study on Goodness of Fit (Jitkrittum et al., 2020), to unlock KSD support on predictive models, we propose to set $P_{\theta}(\mathbf{x})\equiv P_{D}(\mathbf{x})$ such that the identical marginal distribution would not contribute to the discrepancy between the joint distributions $P_{D}(\mathbf{x},y)$ and $P_{\theta}(\mathbf{x},y)$. In addition, while original distribution $P(\mathbf{x})$ that generates data could be arbitrary complex distribution (that out of modelling scope of the predictive model), we may simply set data point distribution $P_{D}$ (not original distribution $P$) as an Uniform distribution over data points in the dataset under the particular context, given a data point is sampled from a generated dataset uniformly. Although the relaxation appears hasty, we believe that the relaxation is valid in the prediction explanation context (and probably only valid in this particular context) in terms of measuring correlation between a pair of data points, where all train/test data points’ input are valid observation (e.g. images) rather than random continuous valued sample in the same space form an unknown distribution.
With the above relaxations, the score function $\nabla_{\mathbf{x},y}\log P_{\theta}(\mathbf{x},y)$ in the Stein operator $\mathcal{A}_{\theta}$ could be derived as a concatenation of the gradient of model $f_{\theta}(\mathbf{x})_{y}$ to its input $\mathbf{x}$ and its probabilistic prediction $f_{\theta}(\mathbf{x})$ in logarithm form, since
@@
$=[\nabla_{\mathbf{x}}\mathbf{y}^{\top}\log f_{\theta}(\mathbf{x})||\nabla_{\mathbf{y}}\mathbf{y}^{\top}\log f_{\theta}(\mathbf{x})]+[\mathbf{0}||\mathbf{0}]$
$=[\nabla_{\mathbf{x}}\log f_{\theta}(\mathbf{x})_{y}||\log f_{\theta}(\mathbf{x})],$
where [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes concatenation operation. We use one-hot vector representation $\mathbf{y}$ to represent data label $y$ here. Since $P_{D}(\mathbf{x})$ follows uniform distribution, its gradient to the inputs is a 0 vector.
In the above derivation, we treat discrete label $y$ as-is without specialized discrete distribution treatments (see (Yang et al., 2018)) to avoid significant computation overhead. In fact, data space in practice is unlikely dense even if a group of features are continuous (e.g. images). Treating the label as a sparse continuous feature can be viewed as an approximation.
@@
Combining Equation 1 and 2, we can estimate the correlation of any pairs of training data points conditioned on the trained machine learning model. Computationally, since a score function $\nabla_{\mathbf{x},y}\log P_{\theta}(\mathbf{x},y)$ depends on a single data point, its outputs of the training set could be pre-computed and cached to accelerate the kernel computation. In particular, the output dimension of the score function is simply $m+k$ for data with $m$ dimensional features and $k$ class labels. Compared to the existing solutions, whose training data cache (or influence) are bounded by the dimension of model parameters (such as Influence function, TracIn, RPS, RPS-JLE), the explanation method built on KSD would come with a significant advantage in terms of scalability (see comparison in Table 1). This statement is generally true for neural network based classifiers, whenever the size of model parameters is far larger than the data dimension.
@@
Figure 2: Demonstration of HD-Explain on 2D Rectangular synthetic dataset. Left shows the training dataset with three classes. Middle figure shows the explanation support of training data points to a given test point (as black cross), where green shows a higher KSD kernel value. Right shows the distribution of KSD kernel values (over the training set) to the test point, where only a small number of training data points provide strong support to this prediction.
3.2 Prediction Explanation
@@
The computation of kernel function in Equation 1 requires access to features and labels of a data point. While the ground-truth label information is available for the training set, it is inaccessible for a test data point. We consider the predicted class $\hat{y}_{t}$ of the test data point $\mathbf{x}_{t}$ as a label to construct a complete data point $(\mathbf{x}_{t},\hat{y}_{t})$ and apply the KSD kernel function. For a test data point $\mathbf{x}_{t}$, we search for top-k training data points that maximize the KSD defined kernel.
@@
Figure 2 demonstrates HD-Explain on a 2d synthetic dataset. The distribution of $\kappa_{\theta}(\mathbf{d},\cdot)$ in the right most plot shows that only a small number of training data points have a strong influence on a particular prediction.
4 Evaluation and Analysis
@@
In this section, we conduct several qualitative and quantitative experiments to demonstrate various properties of HD-Explain and compare it with the existing example-based solutions.
@@
Datasets: We consider multiple disease classification tasks where diagnosis explanation is highly desired. We also introduced synthetic and benchmark classification datasets to deliver the main idea without the need for medical background knowledge. Concretely, we use CIFAR-10 ($32\times 32\times 3$), Brain Tumor (Magnetic Resonance Imaging, $128\times 128\times 3$), Ovarian Cancer (Histopathology Images, $128\times 128\times 3$) datasets, and SVHN ($32\times 32\times 3$). More details are listed in the Appendix G.
@@
Baselines: The baseline explainers used in our experiments include Influence Function, Representer Point Selection, and TracIn. While other variants of these baseline explainers exist (Barshan et al., 2020; Sui et al., 2021; Chen et al., 2021), we note they don’t offer fundamental performance improvements over the classic ones. In addition, as Influence Function and TracIn face scalability issues, we limit the influence of parameters to the last layer of the model so that they can work with models that contain a large number of parameters. Our experiments use ResNet-18 as the backbone model architecture (with around $11$ million trainable parameters) for all image datasets (see Appendix I for detail on our hardware setup). Finally, we also introduce an HD-Explain variant (HD-Explain*) to match the last layer setting of other baseline models, even though HD-Explain can scale up to the whole model without computation pressure (see Appendix J for details).
@@
Metrics: In existing example-based explanation works, the experimental results are often demonstrated qualitatively, as visualized explanation instances, without quantitative evaluation. This results in subjective evaluation. In this paper, we propose several quantitative evaluation metrics to measure the effectiveness of each method.
@@
  • Hit Rate: Hit rate measures how likely an explanation sample hits the desired example cases where the desired examples are guaranteed to be undisputed. Specifically, we modify a training data point with minor augmentations (adding noise or flipping horizontally) and use it as a test data point, such that the best explanation for the generated test data point should be the original data point in the training set.
@@
  • Coverage: Given n test data points, the metric measures the number of unique explanation samples an explanation method produces when configuring to return top-k training samples. Formally,
@@
    where $e_{i}$ is the set of top-k explanations for test data point $i$. Coverage is motivated to measure the diversity of explanations across a test set where a high value reflects higher granularity (per test point) of the explanation.
  • Run Time: It measures the run time of an explanation method in wall clock time.